OpenStudy (anonymous):

what will be the integration of 1/(1+x²)½ along x?

OpenStudy (anonymous):

\[\int\limits 1/(1+x ^{2})^{1/2}\] it will be \[\ln (1 + x ^{2})^{1/2}\]

OpenStudy (anonymous):

you know that when the numerator is the derivative of the dominator then the integration is ln "dominator" :)

OpenStudy (anonymous):

but the derivative of denominator should be there in the numerator

OpenStudy (anonymous):

\[x= \tan \theta\]

OpenStudy (anonymous):

ok , the derivative of (x^2 )^1/2 is 1/2 (2x) so 2 cancel the other 2 :)

OpenStudy (anonymous):

and for natural log, exponent should be 1

OpenStudy (anonymous):

\[\int\limits 1/x= \ln x\]

OpenStudy (anonymous):

if the exponent is other than 1, we use power rule

OpenStudy (anonymous):

yeah but we can't multiply it with 2x so ln(1+x²)½ is not the right answer

OpenStudy (anonymous):

\[d x= \sec^2 \theta d \theta\]

OpenStudy (anonymous):

i think its correct .. anyways good luck ..

OpenStudy (anonymous):

\[\int\limits \sec^2 \theta/(\sqrt{1+\tan^2 \theta} )\]

OpenStudy (anonymous):

\[\int\limits \sec^2\theta/\sec \theta d \theta\]

OpenStudy (anonymous):

\[\int\limits \sec \theta d \theta\]

OpenStudy (anonymous):

\[\ln (\sec \theta + \tan \theta) +C\]

OpenStudy (anonymous):

by backward substitution \[\tan \theta =x\]

OpenStudy (anonymous):

yeah that will give us integration of secθ that is ln(secθ+tanθ) and by substituting and we will get probably get ln(sec(tan^-x)+x)

OpenStudy (anonymous):

\[\sec \theta =\sqrt{1+x^2}\]

OpenStudy (anonymous):

thanks for the help i have some questions on multiple integral usage in finding area and volume. Anyone please help me

OpenStudy (anonymous):

post it as a new Q,some one would surly look at it :)

OpenStudy (anonymous):

Uzma are you from Pakistan?

OpenStudy (anonymous):

ahaa ! ok waiting ur new Qs :)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Can you please tell me your qualification and city

OpenStudy (anonymous):

what would u do with that :)

OpenStudy (anonymous):

nothing just to guess the level of your understanding mathematics

OpenStudy (anonymous):

mathematics is as vast, that one can hardly grasp even one area